Joe is twice as old as David. The sum of their ages is 63. How old are David and Joe
Found 2 solutions by stanbon, JulietG:
Answer by stanbon(75887)

-------
Joe is twice as old as David.
The sum of their ages is 63.
How old are David and Joe
===============
Equations:
j = 2d
j+d = 63
-----
Substitute for "j" and solve for "d":
2d + d = 63
3d = 63
David = 21 years
------------------------
Solve for "j":
j = 2d
Joe = 2*21 = 42 years
------------------------
Cheers,
Stan H.
----------------

Answer by JulietG(1812)

J = 2D
J + D = 63
To solve this, substitute the value of J into the second equation.
J + D = 63
(2D) + D = 63
3D = 63
Divide each side by 3
D = 21
.
If David is 21, then Joe is 42 -- twice as old